A stone is thrown vertically upward with an initial velocity of 40 m/s. Taking g = 10 m/s2, find the maximum height reached by the stone. What is the net displacement and the total distance covered by the stone?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

u = 40 m/s
g = –10 m/s^2(going against gravity)
h = s = ?
v = 0
v^2– u^2= 2gs
(0)^2– (40)^2= 2 (–10) x s
s = –40 x 40 / 2 x –10
s = 80m
Net displacement of the stone = 0 (As the stone falls, back to the same point.)
Total distance covered by stone = 80 m + 80 m
(up)(down)
= 160 m

Thanks